Ryan buys some jumpers to sell on a stall.

He spends £160 buying 60 jumpers.

He sells 80% of the jumpers for £12 each.

He then puts the rest of the jumpers on a Buy one get one half price offer.

He manages to sell half the remaining jumpers using this offer.

How much profit does Ryan make?


1
Expert's answer
2021-12-15T08:44:14-0500

Return from 80% of jumpers sold for 12 each: "60*0.8*12=576."

Remaining jumpers: "60-60*0.8=12."

From these jumpers "12*0.5=6" were sold.

From 6 jumpers 3 were sold for 12 and 3 for "12*0.5=6."

Total return: "576+3*12+3*6=630."

Total profit: "630-160=470."